TCCON > Sites > Eureka > Instrument History > Site Visit 2018-02-24 to 2018-03-15

Site Visit 2018-02-24 to 2018-03-15

Table of contents
No headers

Mar 2

Overcast in the morning, starting to clear in the afternoon but the signal was not strong enough (amplitude < 300 on the InSb in "check signal"), no measurements.

 

On March 1 we did not lock the scanner mechanism correctly, the allen key was "stuck" between the back side of the mechanism and the magnetic voice coil outer instead of just resting on the magnetic voice coil inner (as shown in page 6 of burker_125_align.pdf), we also did not enable the led light with a jumper.

We did this today and took pictures of the fringe pattern at all OPDs with an aperture of 4 mm and 1 mm. With the scanner mechanism correctly locked it looks quite diffrent from yesterday's pictures. We only see a slight shear misalgnment, and a small angular misalignment (at 180 cm OPD and 1mm aperture, we can count 1 more dark fringe at the top than at the bottom).

 

Issues:

We use an external HeNe laser with a beam expander and a biconcave lens to expand the beam. It fills the OAP in the source compartment like the solar beam would. But the laser beam on the aperture wheel is ~5cm in diameter (vs ~5mm for the sun beam).

After the collimating OAP the beam is ~1cm in diameter with the 4 mm aperture, so it does not fill the beamsplitter.

To look at the fringes, we pick off the recombined beam when it is still collimated, with mirrors of diameter ~1 inch.

The equipment we have is not adequate to pick off and view the focused beam between the exit aperture and the last flat mirror.

 

Mar. 1

Overcast skies, no measurements.

We used the alignment kit again to look at the fringes and took pictures at several OPDs. It looked like there was a shear misalignment

 

Feb. 28

Overcast skies, no measurements.

Erik did a N2O cell test.

Both N2O and HBr have the same ME at 180 cm, at ~0.8

We used the kit to look at the fringe pattern, but did not spend enough time to get the pictures.

Before leaving we set up another 1000 scans measurement of the MIR source but with filter 6.

 

Feb. 27

Too windy to operate the dome, no measurements.

I started to put the alignment kit together.

Erik did a HBr cell test.

We set a 1000 scans measurement of the MIR source with filter 3 to check for channeling.

 

Feb. 26

Arrived at ~9 am, today is also cloudy so I ran another HCl test, ME at 45 cm is back at ~1.015 as before the instrument was winterized.

 

Feb. 25

We arrived at the lab at ~1 pm. 

Sébastien cleaned the scanner rods, there was some brown solid deposits on some parts of the rods. Compartment seals were also cleaned and vacuum grease applied on the o-rings.

It was too cloudy to measure, Erik installed the source chiller and Sébastien did a HCl cell test. ME at MOPD is ~ 1.05, up from ~1.015 before the instrument was winterized ...

 

 

Tag page